OpenStudy (anonymous):

Which of the following makes this a true equation? What number plus one and one-fourth equals three and three-fourths? two and one-half two and three-fourths three and one-fourth three and one-half

OpenStudy (anonymous):

@aryandecoolest

OpenStudy (anonymous):

I didn't get the question..can you show.

OpenStudy (anonymous):

@aryandecoolest

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

answer is 5/2

OpenStudy (anonymous):

It's not achoice doe @aryandecoolest

OpenStudy (anonymous):

.dude that's the answer.....2.5=5/2

OpenStudy (anonymous):

Is it a b or c or d? @aryandecoolest

OpenStudy (anonymous):

(A)....2+1/2=5/2

OpenStudy (anonymous):

yes its 5/2